# Closed-loop Uplink Radio Resource Management in CF-O-RAN Empowered 5G Aerial Corridor

### v1

@@ VI Conclusion
This paper presents a joint UAV-O-RU association and power allocation framework for O-RAN-enabled CF mMIMO systems supporting 5G aerial corridors. The key contributions include: (i) an adaptive association algorithm that dynamically selects serving O-RUs by combining UAV and O-RU-centric clustering, followed by QoS-driven refinement for weak UAVs, (ii) a computationally efficient power control method that integrates bisection search with fixed-point iteration achieving global optimality, and (iii) an AO framework integrating both solutions. The framework adopts a CKM within the O-RAN non-RT RIC to mitigate CSI acquisition overhead. Simulation results demonstrated substantial enhancements in minimum SE, QoS compliance, fairness, and computational efficiency compared to baseline approaches. Our analysis reveals that association design is critical for QoS guarantees, while power control dominates fairness. The proposed schemes yield optimal solution within sub-second runtime, suitable for O-RAN complaint real-time deployment. Future work will investigate joint UAV trajectory optimization to enhance coverage and SE in dynamic aerial networks.

### v2

@@ VI Conclusion
This paper presents a joint UAV-O-RU association and power allocation framework for O-RAN-enabled CF mMIMO systems supporting 5G aerial corridors. The key contributions include: (i) an adaptive association algorithm that dynamically selects serving O-RUs by combining UAV and O-RU-centric clustering, followed by QoS-driven refinement for weak UAVs, (ii) a computationally efficient power control method that integrates bisection search with fixed-point iteration achieving global optimality, and (iii) an AO framework integrating both solutions. The framework adopts a CKM within the O-RAN non-RT RIC to mitigate CSI acquisition overhead. Simulation results demonstrated substantial enhancements in minimum SE, QoS compliance, fairness, and computational efficiency compared to baseline approaches. Our analysis reveals that association design is critical for QoS guarantees, while power control dominates fairness. The proposed schemes yield an optimal solution within sub-second runtime, suitable for O-RAN compliant real-time deployment. Future work will investigate joint UAV trajectory optimization to enhance coverage and SE in dynamic aerial networks.